# Towards Synthesized and Editable Motion In-Betweening Through Part-Wise Phase Representation

### v2

@@ Abstract
Styled motion in-betweening is crucial for computer animation and gaming. However, existing methods typically encode motion styles by modeling whole-body motions, often overlooking the representation of individual body parts. This limitation reduces the flexibility of infilled motion, particularly in adjusting the motion styles of specific limbs independently. To overcome this challenge, we propose a novel framework that models motion styles at the body-part level, enhancing both the diversity and controllability of infilled motions. Our approach enables more nuanced and expressive animations by allowing precise modifications to individual limb motions while maintaining overall motion coherence. Leveraging phase-related insights, our framework employs periodic autoencoders to automatically extract the phase of each body part, capturing distinctive local style features. Additionally, we effectively decouple the motion source from synthesis control by integrating motion manifold learning and conditional generation techniques from both image and motion domains. This allows the motion source to generate high-quality motions across various styles, with extracted motion and style features readily available for controlled synthesis in subsequent tasks. Comprehensive evaluations demonstrate that our method achieves superior speed, robust generalization, and effective generation of extended motion sequences. The project page is at https://jixiii.github.io/MIB-BodyPartPhase/.
@@ 1 Introduction
Motion in-betweening is a fundamental technique for generating complete motion sequences from sparse keyframe poses. It serves as a cornerstone for numerous applications in computer vision and computer graphics [41], including character animation [42, 40], action recognition, and motion capture. The diversity and controllability of in-between motion are of paramount importance, as they directly influence the realism, expressiveness, and adaptability of synthesized motion. In this work, we introduce a novel approach to fine-grained style control in motion in-betweening, explicitly focusing on the independent manipulation of motion styles across different body parts.
@@ 3.3 Motion Sampler
The first stage trains the BPMoE using three fundamental components. Reconstruction Loss ($L_{\text{rec}}$): The L2 distance $\|X_{S}-X^{\prime}_{S}\|_{2}^{2}$ between predicted ($X^{\prime}_{S}$) and ground truth ($X_{S}$) joint parameters enforces precise pose reconstruction, preserving local limb geometries and global body coordination. Latent Regularization ($L_{\text{kl}}$): The KL-divergence term $-\frac{1}{2}(1+\sigma^{2}-\mu^{2}-e^{\sigma})$ regularizes the control signal distribution $z\sim\mathcal{N}(0,1)$, preventing overfitting while maintaining a smooth latent space for diverse motion sampling. The hyperparameter $\beta=0.001$ balances reconstruction accuracy and generative diversity [32]. Foot Plant Constraint ($L_{\text{foot}}$): To eliminate foot sliding, we introduce $\|f^{\prime}_{v}\odot\delta(v_{f})\|_{2}^{2}$ where $\delta(v_{f})$ implements velocity-dependent contact weighting. The composite BPMoE objective integrates these components:
@@
, where phase velocity $F^{t}$ preserves cyclic continuity via: $F^{t}=(S^{t}-S^{t-1})-\lfloor S^{t}-S^{t-1}+0.5\rfloor$The sampler’s total loss combines these objectives:
@@
Frames 40 80 120 140 160 40 80 120 140 160 40 80 120 140 160
CVAE[32] 0.120 1.837 7.179 10.788 14.645 3.208 13.090 27.936 34.322 39.731 0.135 0.164 0.127 0.111 0.097
RSMT[33] 0.068 2.146 7.378 10.311 13.981 2.179 11.828 25.848 32.082 36.664 0.043 0.131 0.137 0.129 0.116
PhaseMIB[28] 0.095 1.870 7.133 10.512 12.171 3.125 13.491 25.768 31.586 35.798 0.104 0.177 0.142 0.132 0.126
OURS(2) 0.083 2.010 6.780 9.443 12.178 2.812 13.536 25.300 30.052 35.185 0.081 0.145 0.116 0.100 0.093
OURS(5) 0.122 2.663 7.800 10.909 14.626 4.097 14.100 25.816 31.118 35.746 0.096 0.131 0.094 0.083 0.074
@@ 4 Experiments
Frames 40 80 120 140 160 40 80 120 140 160 40 80 120 140 160
CVAE[32] 0.157 1.869 7.279 10.968 14.821 0.171 1.876 7.289 10.852 14.763 0.171 1.945 7.253 10.901 14.817
RSMT(original)[33] 0.109 2.261 7.266 10.525 14.113 0.100 2.234 7.312 10.567 14.038 0.101 2.223 7.384 10.453 14.178
RSMT(modified)[33] 0.111 2.191 7.297 10.426 13.909 0.107 2.175 7.221 10.466 13.918 0.108 2.162 7.372 10.521 14.197
PhaseMIB[28] 0.116 1.876 7.158 10.524 12.985 0.112 2.143 7.667 10.686 12.769 0.120 2.183 7.731 10.705 12.810
OURS(2) 0.114 2.147 6.938 9.633 12.364 0.108 2.148 6.928 9.603 12.333 0.106 2.194 6.885 9.634 12.307
OURS(5) 0.139 2.719 7.779 10.943 14.710 0.139 2.707 7.795 10.950 14.455 0.137 2.717 7.913 11.075 14.672
@@
CVAE[32] 4.105 13.515 28.343 34.586 40.027 4.230 13.510 28.220 34.440 40.036 4.228 13.537 28.256 34.621 39.970
RSMT(original)[33] 3.000 12.289 26.314 32.333 36.925 2.861 12.187 26.087 32.196 36.899 2.876 12.217 26.233 32.297 37.042
RSMT(modified)[33] 3.011 12.229 26.255 32.303 36.998 2.997 12.225 26.074 32.164 36.790 3.013 12.240 26.185 32.219 37.033
PhaseMIB[28] 3.211 13.653 27.124 31.714 35.946 3.614 13.703 26.537 33.322 36.754 3.769 13.857 27.934 34.386 36.967
OURS(2) 3.511 14.134 26.027 30.781 35.701 3.353 13.926 25.798 30.518 35.387 3.316 13.929 25.675 30.451 35.454
OURS(5) 4.512 14.456 26.113 31.543 36.242 4.451 14.354 25.995 31.424 36.029 4.461 14.454 26.162 31.342 35.969
@@
GT 0.045 0.044 0.044 0.044 0.044 0.045 0.044 0.044 0.044 0.044 0.045 0.044 0.044 0.044 0.044
[…]
RSMT(original)[33] 0.046 0.131 0.141 0.134 0.120 0.045 0.134 0.141 0.134 0.122 0.045 0.131 0.138 0.130 0.117
RSMT(modified)[33] 0.047 0.131 0.141 0.131 0.119 0.045 0.132 0.142 0.130 0.120 0.045 0.131 0.138 0.129 0.118
PhaseMIB[28] 0.097 0.133 0.146 0.139 0.126 0.098 0.133 0.152 0.137 0.130 0.102 0.135 0.154 0.143 0.137
OURS(2) 0.083 0.149 0.115 0.105 0.094 0.081 0.146 0.115 0.102 0.091 0.084 0.147 0.116 0.105 0.093
OURS(5) 0.099 0.135 0.099 0.085 0.078 0.099 0.133 0.095 0.083 0.076 0.096 0.134 0.099 0.086 0.077
@@
Frames 40 80 120 140 160 40 80 120 140 160 40 80 120 140 160
Motion-based 0.115 2.173 7.573 10.927 14.616 4.248 14.963 27.551 32.609 37.184 0.126 0.186 0.148 0.126 0.109
OURS(2) 0.083 2.010 6.780 9.443 12.178 2.812 13.536 25.300 30.052 35.185 0.081 0.145 0.116 0.100 0.093
OURS(5) 0.122 2.663 7.800 10.909 14.626 4.097 14.100 25.816 31.118 35.746 0.096 0.131 0.094 0.083 0.074
Table 3: Comparisons of reconstruction and foot skating between motion-based representation and body-part-phase-based representation.
@@ 4.1 Experimental Settings
To accelerate training convergence, we randomly select a 25-frame window from the 120-frame clips for each step instead of using the full sequence. We employ AMSGrad with parameters $(\beta_{1}=0.5,\beta_{2}=0.9)$ and a learning rate of $31e-$ 3. For training the sampler, we shuffle all style sequences randomly. To ensure model robustness across varying time lengths, we randomly sample a sequence starting at a length of 20 frames and linearly increase it to 40 frames at each epoch, similar to [7]. We use the same AMSGrad setup as before, with a weight decay of $41e-$ 4 for training the style encoder to prevent overfitting, while the weight decay for other modules is set to 0. After training the BPMoE, we remove its encoder, fix its decoder, and connect the motion sampler to the fixed decoder to train the sampler. This training process typically takes around a day. We conduct our experiments using an NVIDIA 4090 24G GPU. Our method takes, on average, 1.87 ms to synthesize one frame, which is sufficient for real-time applications.

### v3

@@ Abstract
Styled motion in-betweening is crucial for computer animation and gaming. However, existing methods typically encode motion styles by modeling whole-body motions, often overlooking the representation of individual body parts. This limitation reduces the flexibility of infilled motion, particularly in adjusting the motion styles of specific limbs independently. To overcome this challenge, we propose a novel framework that models motion styles at the body-part level, enhancing both the diversity and controllability of infilled motions. Our approach enables more nuanced and expressive animations by allowing precise modifications to individual limb motions while maintaining overall motion coherence. Leveraging phase-related insights, our framework employs periodic autoencoders to automatically extract the phase of each body part, capturing distinctive local style features. Additionally, we effectively decouple the motion source from synthesis control by integrating motion manifold learning and conditional generation techniques from both image and motion domains. This allows the motion source to generate high-quality motions across various styles, with extracted motion and style features readily available for controlled synthesis in subsequent tasks. Comprehensive evaluations demonstrate that our method achieves superior speed, robust generalization, and effective generation of extended motion sequences.
@@ 1 Introduction
Motion in-betweening is a fundamental technique for generating complete motion sequences from sparse keyframe poses. It serves as a cornerstone for numerous applications in computer vision and computer graphics, including character animation [41, 40], action recognition, and motion capture. The diversity and controllability of in-between motion are of paramount importance, as they directly influence the realism, expressiveness, and adaptability of synthesized motion. In this work, we introduce a novel approach to fine-grained style control in motion in-betweening, explicitly focusing on the independent manipulation of motion styles across different body parts.
@@ 3.3 Motion Sampler
The first stage trains the BPMoE using three fundamental components. Reconstruction Loss ($L_{\text{rec}}$): The L2 distance $\|X_{S}-X^{\prime}_{S}\|_{2}^{2}$ between predicted ($X^{\prime}_{S}$) and ground truth ($X_{S}$) joint parameters enforces precise pose reconstruction, preserving local limb geometries and global body coordination. Latent Regularization ($L_{\text{kl}}$): The KL-divergence term $\frac{1}{2}(1+\sigma^{2}-\mu^{2}-e^{\sigma})$ regularizes the control signal distribution $z\sim\mathcal{N}(0,1)$, preventing overfitting while maintaining a smooth latent space for diverse motion sampling. The hyperparameter $\beta=0.001$ balances reconstruction accuracy and generative diversity [32]. Foot Plant Constraint ($L_{\text{foot}}$): To eliminate foot sliding, we introduce $\|f^{\prime}_{v}\odot\delta(v_{f})\|_{2}^{2}$ where $\delta(v_{f})$ implements velocity-dependent contact weighting. The composite BPMoE objective integrates these components:
@@
, where phase velocity $F^{t}$ preserves cyclic continuity via: $F^{t}=(S^{t}-S^{t-1})-\lfloor S^{t}-S^{t-1}+0.5\rfloor$ .The sampler’s total loss combines these objectives:
@@
Frames 120 140 160 120 140 160 120 140 160
CVAE 7.179 10.788 14.645 27.936 34.322 39.731 0.127 0.111 0.097
RSMT 7.378 10.311 13.981 25.848 32.082 36.664 0.137 0.129 0.116
PhaseMIB 7.133 10.512 12.171 25.768 31.586 35.798 0.142 0.132 0.126
OURS(2) 6.780 9.443 12.178 25.300 30.052 35.185 0.116 0.100 0.093
OURS(5) 7.800 10.909 14.626 25.816 31.118 35.746 0.094 0.083 0.074
@@ 4 Experiments
Frames 120 140 160 120 140 160 120 140 160
CVAE 7.279 10.968 14.821 7.289 10.852 14.763 7.253 10.901 14.817
RSMT(original) 7.266 10.525 14.113 7.312 10.567 14.038 7.384 10.453 14.178
RSMT(modified) 7.297 10.426 13.909 7.221 10.466 13.918 7.372 10.521 14.197
PhaseMIB 7.158 10.524 12.985 7.667 10.686 12.769 7.731 10.705 12.810
OURS(2) 6.938 9.633 12.364 6.928 9.603 12.333 6.885 9.634 12.307
OURS(5) 7.779 10.943 14.710 7.795 10.950 14.455 7.913 11.075 14.672
@@
120 140 160 120 140 160 120 140 160
CVAE 28.343 34.586 40.027 28.220 34.440 40.036 28.256 34.621 39.970
RSMT(original) 26.314 32.333 36.925 26.087 32.196 36.899 26.233 32.297 37.042
RSMT(modified) 26.255 32.303 36.998 26.074 32.164 36.790 26.185 32.219 37.033
PhaseMIB 27.124 31.714 35.946 26.537 33.322 36.754 27.934 34.386 36.967
OURS(2) 26.027 30.781 35.701 25.798 30.518 35.387 25.675 30.451 35.454
OURS(5) 26.113 31.543 36.242 25.995 31.424 36.029 26.162 31.342 35.969
@@
120 140 160 120 140 160 120 140 160
GT 0.044 0.044 0.044 0.044 0.044 0.044 0.044 0.044 0.044
CVAE 0.132 0.114 0.100 0.130 0.112 0.099 0.130 0.113 0.100
RSMT(original) 0.141 0.134 0.120 0.141 0.134 0.122 0.138 0.130 0.117
RSMT(modified) 0.141 0.131 0.119 0.142 0.130 0.120 0.138 0.129 0.118
PhaseMIB 0.146 0.139 0.126 0.152 0.137 0.130 0.154 0.143 0.137
OURS(2) 0.115 0.105 0.094 0.115 0.102 0.091 0.116 0.105 0.093
OURS(5) 0.099 0.085 0.078 0.095 0.083 0.076 0.099 0.086 0.077
@@
Frames 120 140 160 120 140 160 120 140 160
Motion-based 7.573 10.927 14.616 27.551 32.609 37.184 0.148 0.126 0.109
OURS(2) 6.780 9.443 12.178 25.300 30.052 35.185 0.116 0.100 0.093
OURS(5) 7.800 10.909 14.626 25.816 31.118 35.746 0.094 0.083 0.074
Table 3: Comparisons of reconstruction and foot skating between motion-based and body-part-phase-based representation.
@@ 4.1 Experimental Settings
To accelerate training convergence, we randomly select a 25-frame window from the 120-frame clips for each step instead of using the full sequence. We employ AMSGrad with parameters $(\beta_{1}=0.5,\beta_{2}=0.9)$ and a learning rate of 1e31e-3. For training the sampler, we shuffle all style sequences randomly. To ensure model robustness across varying time lengths, we randomly sample a sequence starting at a length of 20 frames and linearly increase it to 40 frames at each epoch, similar to [7]. We use the same AMSGrad setup as before, with a weight decay of 1e41e-4 for training the style encoder to prevent overfitting, while the weight decay for other modules is set to 0. After training the BPMoE, we remove its encoder, fix its decoder, and connect the motion sampler to the fixed decoder to train the sampler. This training process typically takes around a day. We conduct our experiments using an NVIDIA 4090 24G GPU. Our method takes, on average, 1.87 ms to synthesize one frame, which is sufficient for real-time applications.